[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Aug 17, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Aug 17 20:22:05 UTC 2017


HSU Shift Log for Thu Aug 17 2017 16:21:58 GMT-0400 (EDT)

Shift: H

Forecaster: J Beven

Technical Issues Summary:

ATCF issues: None

N-AWIPS issues: Lost a monitor on Ridges around 2:00pm due to one of the  
extenders becoming unplugged. DAZ and Brian were able to fix it.

AWIPS-2/NCP issues: None

Other Techincal issues: None

Miscellaneous issues: GOES-16 meso 1-minute imagery was very useful in  
finding the center of Nine/Harvey this morning and throughout the day.  
Barbados Radar down until September.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170817/b983684d/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list